OpenStudy (anonymous):

]What is the simplified form of the quotient of 3 x and the sum of 4 x and 8 all over the quotient of 7 x and the expression x squared minus 4 x minus 12?

OpenStudy (anonymous):

3x(4x+8) ____________ 7x/x^2-4x-12 solve that.

OpenStudy (kropot72):

The question can be rearranged as \[\frac{3x}{4x+8}\times \frac{x ^{2}-4x-12}{7x}\] The denominator of the first expression and the numerator of the second expression can be factorised. Then you can simplify.